OpenStudy (anonymous):

Multiply and simplify: (I have to type the equation again in equation form otherwise it will be confusing.)

OpenStudy (anonymous):

\[\sqrt{3}(4+\sqrt{12})\]

OpenStudy (anonymous):

Answer choices: A. \[\sqrt{12}+6\] B. \[4\sqrt{3}+6\] C. \[2\sqrt{3}+6\sqrt{3}\]

OpenStudy (anonymous):

B

OpenStudy (anonymous):

The answer will be B :)

OpenStudy (lalaly):

\[\sqrt{12}=\sqrt4 \times \sqrt3= 2\sqrt3\]\[4\sqrt3+ (2 \times 3) = 6+4\sqrt3\]